# BeetleFlow: An Integrative Deep Learning Pipeline for Beetle Image Processing

### v1

@@ 4.1.1 Beetle Detection
We apply the detection pipeline on the dataset collected by the National Ecological Observatory Network (NEON) from ecological sites across the U.S., along with associated metadata. The dataset contains 1,506 images of trays containing pinned carabid specimens. The beetles were imaged following the optimized image capture guidelines for biodiversity specimens [25].
@@ 4.1.2 Beetle Segmentation
The unlabeled individual beetle images are derived from our pipeline by processing the tray images. For 5-class labeling, we manually labeled 160 beetles and utilized an additional 180 labeled beetles from a previous work, SST [26]. A total of 340 labeled beetles were then partitioned into a training set of 272 and a test set of 68 images. For 9-class labeling, we manually labeled 330 beetles, dividing them into a training set of 264 and a test set of 66 images.

### v2

@@ 4.1.1 Beetle Detection
We apply the detection pipeline on the dataset collected by the National Ecological Observatory Network (NEON) [20, 21] from ecological sites across the U.S., along with associated metadata. The dataset contains 1,506 images of trays containing pinned carabid specimens. The beetles were imaged following the optimized image capture guidelines for biodiversity specimens [6].
@@ 4.1.2 Beetle Segmentation
The unlabeled individual beetle images are derived from our pipeline by processing the tray images. For 5-class labeling, we manually labeled 160 beetles and utilized an additional 180 labeled beetles from a previous work, SST [7, 8, 23]. A total of 340 labeled beetles were then partitioned into a training set of 272 and a test set of 68 images. For 9-class labeling, we manually labeled 330 beetles, dividing them into a training set of 264 and a test set of 66 images.
@@ 5 Discussion and Future Work
Acknowledgments and Disclosure of Funding
This work is supported by the National Science Foundation OAC 2118240 Imageomics Institute award and is based in part upon work supported by the National Ecological Observatory Network (NEON), a program sponsored by the U.S. National Science Foundation (NSF) and operated under cooperative agreement by Battelle. The authors are grateful for the generous support of the computational resources from the Ohio Supercomputer Center.